# SeaS: Few-shot Industrial Anomaly Image Generation with Separation and Sharing Fine-tuning

### v1

@@ A.4 Additional dataset results
Category DFMGAN (Duan et al., 2023) AnomalyDiffusion (Hu et al., 2024) Ours
@@
Figure 7: Visualization of the generation results on MVTec 3D AD. The sub-image in the lower right corner is the generated mask.
Comparison Methods
In terms of compared approaches, since existing state-of-the-art approaches, e.g., DFMGAN(Duan et al., 2023), AnomalyDiffusion (Hu et al., 2024), conducted the experimental evaluations only on MVTec AD dataset (Bergmann et al., 2019), we evaluate them on MVTec 3D AD dataset (Bergmann. et al., 2022) using their official source codes.
Anomaly image generation quality
As presented in Tab. 8, SeaS achieves scores of 1.95 on IS and 0.30 on IC-LPIPS, demonstrating our method’s ability to generate anomaly images with superior fidelity and diversity. The generated anomaly images are shown in Fig. 7. SeaS can generate images with diverse anomalies, avoiding overfitting to the training images (e.g., cookie_crack and foam_color), while ensuring the fidelity of the generated images (e.g., cable_gland_bent). Additionally, the masks generated by our method are accurately aligned with the anomalies (e.g., peach_cut).
Anomaly segmentation and detection
Tab. 9 shows the comparisons on downstream supervised segmentation trained by the generated images. It consistently demonstrates that our method outperforms others across all the segmentation models, with a 15.49% average improvement on IoU. The segmentation anomaly maps are shown in Fig. 8. There are fewer false positives (e.g., potato_combined) and fewer false negatives (e.g., bagel_contamination), when the BiSeNet V2 is trained on the image-mask pairs generated by our method. In addition, we report the image-level metrics in Tab. 10, and our method achieves a 6.74% gain on image-AUROC. Tab. 11 shows the comparisons of anomaly detection methods HVQ-Trans (Lu et al., 2023), Shape-guided (Chu et al., 2023), and FOD (Yao et al., 2023) on anomaly segmentation tasks. Supervised segmentation models achieve better performance than most unsupervised AD methods on small-scale networks, with an IoU of 39.00% on LFD (0.936M). The pixel-level AUROC, which is sensitive to false negatives but less sensitive to false positives, of the Shape-guided method is higher. However, our observation indicates that the Shape-guided method has a high number of false positives. This significantly degrades the segmentation metrics, resulting in low pixel-level AP, $F_{1}$-max, and IoU scores. For industrial anomaly detection, an effective method should achieve a balance between false positives and false negatives.
Figure 8: Qualitative anomaly segmentation results with BiSeNet V2 on MVTec 3D AD.
Table 9: Comparison on anomaly segmentation on MVTec 3D AD.
Model DFMGAN (Duan et al., 2023) AnomalyDiffusion (Hu et al., 2024) Ours
AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU
BiSeNet V2 (Yu et al., 2021) 75.89 15.02 21.73 15.68 92.39 15.15 20.09 14.70 90.41 26.04 32.61 28.55
UPerNet (Xiao et al., 2018) 75.12 19.54 26.04 18.78 88.48 28.95 35.81 25.04 91.93 38.51 43.53 38.56
LFD (Zhou et al., 2024a) 72.15 9.54 14.29 14.81 92.68 24.29 32.74 19.90 91.61 40.25 43.47 39.00
Average 74.39 14.70 20.69 16.42 91.18 22.80 29.55 19.88 91.32 34.93 39.87 35.37
Table 10: Comparison on image-level anomaly detection on MVTec 3D AD.
Model DFMGAN (Duan et al., 2023) AnomalyDiffusion (Hu et al., 2024) Ours
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max
BiSeNet V2 (Yu et al., 2021) 61.88 81.80 84.44 61.49 81.35 85.36 73.60 87.75 85.82
UPerNet (Xiao et al., 2018) 67.56 84.53 84.99 76.56 90.42 87.35 82.75 92.59 88.72
LFD (Zhou et al., 2024a) 62.23 82.17 85.38 77.06 89.44 87.20 78.96 91.22 87.28
Average 63.89 82.83 84.94 71.70 87.07 86.64 78.44 90.52 87.27
Table 11: Comparison with anomaly detection methods on MVTec 3D AD.
Model Parameters Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU
SeaS + BiSeNet V2 (Yu et al., 2021) 3.341M 73.60 87.75 85.82 90.41 26.04 32.61 28.55
SeaS + UPerNet (Xiao et al., 2018) 64.042M 82.57 92.59 88.72 91.93 38.51 43.53 38.56
SeaS + LFD (Zhou et al., 2024a) 0.936M 78.96 91.22 87.28 91.61 40.25 43.47 39.00
HVQ-Trans (Lu et al., 2023) 8.45M 68.15 84.38 85.20 96.40 24.59 17.23 20.51
Shape-guided (Chu et al., 2023) 4.13M 79.07 91.05 88.72 98.45 26.69 34.16 34.12
FOD (Yao et al., 2023) 3.58M 71.66 86.83 86.57 97.03 14.70 20.99 23.31
More qualitative anomaly image generation results
We provide further qualitative results of every category on the MVTec 3D AD dataset, from Fig. 9 to Fig. 10. We report the anomaly image generation results of SeaS for varying types of anomalies. The first column represents the generated anomaly images, the second column represents the corresponding generated masks.
Figure 9: Qualitative results of our anomaly image generation results on MVTec 3D AD. In the first row, from left to right are the results for bagel, cable_gland, and carrot categories. In the second row, from left to right are the results for cookie, dowel, and foam categories.
Figure 10: Qualitative results of our anomaly image generation results on MVTec 3D AD. In the first row, from left to right are the results for peach, and potato categories. In the second row, from left to right are the results for rope, and tire categories.
More qualitative segmentation results with different segmentation models
In this section, we provide further qualitative results with the anomaly segmentation models on the MVTec 3D AD dataset. As shown in Fig. 11, we report the segmentation results of different segmentation models trained on image-mask pairs generated by SeaS.
Figure 11: Qualitative comparison results with the anomaly segmentation models on MVTec 3D AD. In the figure, from top to bottom are the results for cable_gland, carrot, dowel, foam and peach categories.
A.5 More quantitative comparison with unsupervised anomaly detection methods
We compare our method with the state-of-the-art unsupervised anomaly detection methods, i.e., RealNet (Zhang et al., 2024), and HVQ-Trans (Lu et al., 2023). The performance is different from the results reported in the paper, since as we mentioned in Sec 4.1, we use 2/3 anomaly images and all good images in the testing set of MVTec AD as the testing set in all the experiments, while the original results are achieved on the whole testing set. As shown in Tab. 12, Real-Net contains 591M parameters, around 177 times larger than BiSeNet V2, and 631 times larger than LFD, while the pixel-level AP and IoU measures of Real-Net are even worse than those of BiSeNet V2 and LFD. Although the pixel-level AUROC metric, which is more sensitive to false negatives than to false positives, is slightly higher for Real-Net, we observe that it generates a high number of false positives, substantially reducing pixel-level AP, $F_{1}$-max, and IoU scores. For effective industrial anomaly detection, a method must balance false positives and false negatives. The supervised anomaly segmentation models greatly outperform HVQ-Trans (8.45M parameters) in AP, F1-max measure, and IoU, even though BiSeNet V2 and LFD are much smaller. These comparisons reveal that the small segmentation model achieves good performance using the generated images of SeaS, which is important for practical industrial applications.
Table 12: Comparison with anomaly detection methods on MVTec AD.
Model Parameters Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU
SeaS + BiSeNet V2 (Yu et al., 2021) 3.341M 96.00 98.14 95.43 97.21 69.21 66.37 55.28
SeaS + UPerNet (Xiao et al., 2018) 64.042M 98.29 99.20 97.34 97.87 74.42 70.70 61.24
SeaS + LFD (Zhou et al., 2024a) 0.936M 95.88 97.89 95.15 98.09 77.15 72.52 56.47
Real-Net (Zhang et al., 2024) 591M 98.19 98.99 97.88 98.84 68.09 66.46 53.99
HVQ-Trans (Lu et al., 2023) 8.45M 96.38 98.09 95.30 97.60 47.95 53.32 45.03
@@ A.6 More qualitative and quantitative anomaly image generation results
Figure 12: Qualitative results of our anomaly image generation results on MVTec AD. In the first row, from left to right are the results for transistor, wood, and zipper categories.
@@
Figure 13: Qualitative results of our anomaly image generation results on MVTec AD. In the first row, from left to right are the results for bottle, cable, and capsule categories. In the second row, from left to right are the results for carpet, grid, and hazelnut categories.
@@
Figure 14: Qualitative results of our anomaly image generation results on MVTec AD. In the first row, from left to right are the results for leather, metal_ nut, and pill categories. In the second row, from left to right are the results for screw, tile, and toothbrush categories.
More quantitative results
@@
In this section, we report the detailed generation results of SeaS for each category on the MVTec AD datasets, compared with DRAEM (Zavrtanik et al., 2021), DFMGAN (Duan et al., 2023) and AnomalyDiffusion (Hu et al., 2024) which are presented from Tab. 13 to Tab. 18
Table 13: Comparison on anomaly segmentation on BiSeNet V2.
Category DRAEM DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU
bottle 77.42 45.25 48.21 45.77 89.34 64.67 62.78 44.71 99.00 88.02 80.53 68.25 99.46 93.43 85.59 75.86
cable 65.28 14.73 23.09 19.44 93.87 67.98 64.74 44.02 92.84 69.86 66.32 46.49 89.85 72.07 71.58 53.24
capsule 63.71 13.31 20.00 37.88 74.88 16.43 23.01 29.97 92.71 38.11 40.67 19.44 86.33 24.64 30.54 39.70
carpet 97.27 69.99 68.28 54.31 94.53 42.53 47.44 39.88 98.65 73.10 65.83 43.25 99.61 82.30 72.94 55.52
grid 93.86 25.36 35.46 35.57 96.86 24.40 37.40 29.93 80.59 8.08 16.79 14.26 99.36 37.91 42.50 39.80
hazelnut 77.48 41.52 48.36 54.58 99.87 96.75 90.07 71.68 97.71 63.34 59.87 43.12 97.82 78.55 73.09 68.47
leather 99.76 67.02 64.96 52.37 97.50 51.10 52.26 50.67 99.30 57.49 59.62 43.94 98.91 59.84 58.62 45.82
metal_nut 73.26 32.26 32.77 48.68 99.39 97.59 92.52 70.40 99.03 95.67 88.69 58.8 99.69 98.29 93.23 74.40
pill 60.02 9.33 17.17 11.67 97.09 83.98 79.26 36.39 99.44 93.16 86.62 41.18 98.31 76.97 68.00 55.43
screw 82.23 17.78 24.08 22.15 97.94 37.10 41.01 31.63 94.08 17.95 25.90 20.00 97.64 40.20 45.35 38.43
tile 98.09 82.9 76.43 63.48 99.65 97.08 91.16 75.94 97.79 85.58 78.28 60.46 99.67 97.29 91.48 75.75
toothbrush 92.65 36.73 45.92 23.90 97.70 51.32 54.05 23.38 98.43 49.64 54.08 26.53 97.15 46.09 49.02 28.56
transistor 62.48 14.83 20.57 21.85 84.31 45.34 46.07 30.00 98.85 85.27 77.95 49.83 96.75 69.52 66.11 57.24
wood 92.89 70.82 68.34 58.05 98.32 64.82 63.11 58.99 96.78 63.38 60.31 45.73 98.38 80.81 74.03 56.22
zipper 84.18 41.68 45.65 41.08 97.29 65.18 63.24 49.93 98.81 78.89 72.66 62.03 99.23 80.27 73.41 64.80
Average 81.37 38.90 42.62 39.39 94.57 60.42 60.54 45.83 96.27 64.5 62.27 42.89 97.21 69.21 66.37 55.28
@@
Category DRAEM DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max
bottle 95.93 98.21 93.18 96.74 98.75 95.35 98.14 99.34 97.67 100.00 100.00 100.00
cable 80.79 83.53 79.22 79.47 85.00 74.13 95.37 96.71 92.91 94.61 96.39 89.83
capsule 91.88 97.62 92.41 85.51 95.16 89.82 84.06 95.01 89.74 88.81 96.92 89.21
carpet 98.21 99.24 95.31 91.42 96.29 88.89 90.55 96.41 90.32 98.16 99.31 97.56
grid 96.43 98.50 95.00 99.64 99.82 97.56 81.19 89.92 83.95 99.17 99.63 98.73
hazelnut 97.92 98.65 94.62 100.00 100.00 100.00 93.39 95.74 90.91 100.00 100.00 100.00
leather 100.00 100.00 100.00 98.31 99.23 95.24 100.00 100.00 100.00 95.83 98.38 95.93
metal_nut 96.38 99.00 96.83 97.37 99.16 94.66 99.01 99.66 97.71 100.00 100.00 100.00
pill 74.68 89.88 89.00 84.86 95.27 91.00 90.38 97.43 91.35 96.59 99.12 95.24
screw 71.15 83.52 83.15 74.95 85.50 80.72 58.18 75.32 81.25 77.24 89.55 80.60
tile 99.68 99.82 98.28 99.47 99.74 99.12 98.78 99.44 97.39 100.00 100.00 100.00
toothbrush 82.50 90.00 85.11 78.33 87.73 83.72 78.33 89.26 79.17 90.42 94.49 89.47
transistor 73.87 68.65 60.87 79.52 75.77 69.57 94.40 94.68 94.34 99.23 98.39 94.92
wood 97.24 98.99 96.30 98.87 99.46 97.67 90.48 94.12 93.33 100.00 100.00 100.00
zipper 91.31 97.01 90.24 98.97 99.64 97.56 98.89 99.62 97.56 100.00 100.00 100.00
Average 89.87 93.51 89.97 90.90 94.43 90.33 90.08 94.84 91.84 96.00 98.14 95.43
Table 15: Comparison on anomaly segmentation on UPerNet.
Category DRAEM DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU
bottle 82.87 51.45 52.25 54.58 87.94 56.89 56.56 45.41 99.54 93.01 85.94 75.31 99.28 91.73 84.53 78.73
cable 56.64 13.59 22.42 16.29 87.52 64.30 65.61 41.02 91.00 68.12 67.49 51.84 91.08 76.25 74.63 59.00
capsule 60.95 11.12 18.34 27.18 67.92 12.31 20.32 30.47 97.64 51.90 51.66 37.00 92.09 39.60 43.89 50.18
carpet 98.83 79.57 72.53 61.50 95.85 36.05 34.52 48.10 99.45 82.13 72.55 53.17 99.67 82.01 73.53 60.60
grid 95.66 36.69 43.55 36.52 97.49 29.67 36.15 31.37 94.22 28.97 38.50 32.93 99.18 44.94 48.28 44.21
hazelnut 77.69 41.57 47.13 58.88 99.36 79.76 71.10 72.90 97.77 70.48 67.93 54.47 99.54 81.84 75.48 73.30
leather 99.70 65.49 63.01 62.20 80.97 17.60 26.21 30.17 99.48 63.46 60.54 48.70 99.42 68.26 65.52 57.01
metal_nut 65.37 23.26 27.39 42.56 98.44 95.64 91.48 64.92 98.62 95.11 88.62 61.31 99.70 98.33 92.90 76.07
pill 64.46 11.33 20.28 13.10 97.58 83.74 80.02 42.33 99.33 95.04 88.77 49.18 98.59 81.16 74.26 62.62
screw 90.88 23.64 31.49 24.71 97.49 53.83 53.02 42.05 93.89 36.60 42.68 34.08 98.97 52.02 51.65 46.61
tile 96.25 79.31 74.18 66.79 99.79 97.29 91.11 77.46 94.70 73.34 67.79 58.54 99.67 95.89 90.71 77.89
toothbrush 93.86 46.93 58.92 26.76 97.42 51.09 59.23 28.33 97.52 60.67 59.46 33.98 98.50 63.62 63.07 42.09
transistor 78.20 26.52 30.34 26.07 82.07 36.31 39.48 27.44 94.26 73.68 69.50 53.64 93.88 70.37 68.12 56.98
wood 95.03 77.07 74.07 64.67 97.90 69.02 62.21 63.10 96.09 70.10 64.38 51.44 99.28 85.28 76.28 65.09
zipper 91.74 54.11 53.69 48.57 97.28 71.60 66.64 54.54 99.54 86.18 78.50 66.47 99.17 85.01 77.57 68.21
Average 83.21 42.78 45.97 42.03 92.33 57.01 56.91 46.64 96.87 69.92 66.95 50.80 97.87 74.42 70.70 61.24
@@
Category DRAEM DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max
bottle 98.49 99.41 97.62 94.19 97.86 93.18 100.00 100.00 100.00 100.00 100.00 100.00
cable 73.06 77.88 71.94 85.64 90.03 80.33 95.58 97.06 92.56 94.40 96.38 92.44
capsule 85.62 95.66 89.02 81.04 94.26 87.01 96.00 98.77 95.48 94.43 98.44 92.21
carpet 97.64 99.00 95.08 96.72 98.58 93.75 98.68 99.53 98.36 99.94 99.97 99.20
grid 97.62 98.99 97.44 98.33 99.13 96.30 96.67 98.73 97.44 99.76 99.88 98.73
hazelnut 97.14 97.74 92.63 99.84 99.87 97.96 99.17 99.43 97.87 100.00 100.00 100.00
leather 100.00 100.00 100.00 79.91 90.70 81.75 100.00 100.00 100.00 100.00 100.00 100.00
metal_nut 86.79 96.02 87.39 98.30 99.38 97.71 98.65 99.62 98.41 99.72 99.91 99.21
pill 67.07 87.37 88.07 88.54 96.56 92.39 91.23 97.78 90.91 98.28 99.58 97.92
screw 80.04 90.95 81.03 89.01 94.54 88.24 85.06 93.87 85.33 93.47 97.07 90.45
tile 99.20 99.51 98.25 99.68 99.81 99.13 99.68 99.81 99.13 100.00 100.00 100.00
toothbrush 86.67 94.19 88.89 75.00 86.99 80.00 90.00 95.13 90.00 95.00 97.65 94.74
transistor 79.29 74.68 69.09 83.04 73.59 74.19 100.00 100.00 100.00 99.52 99.16 96.43
wood 98.25 99.29 96.47 93.36 95.60 95.45 98.62 99.49 97.62 99.87 99.94 98.82
zipper 94.86 98.13 92.02 98.48 99.51 98.14 100.00 100.00 100.00 100.00 100.00 100.00
Average 89.45 93.92 89.66 90.74 94.43 90.37 96.62 98.61 96.21 98.29 99.20 97.34
Table 17: Comparison on anomaly segmentation on LFD.
Category DRAEM DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU
bottle 78.76 50.40 51.05 39.74 90.41 61.51 58.49 40.19 98.71 89.64 81.55 67.10 99.28 92.65 84.86 73.82
cable 67.64 19.41 25.36 19.19 96.49 79.40 75.25 53.47 97.89 79.85 72.75 53.69 94.53 75.41 72.70 55.98
capsule 88.48 34.60 39.62 31.48 91.82 56.11 58.56 32.50 95.80 38.17 48.92 32.04 91.80 49.76 53.69 41.14
carpet 83.51 37.69 41.86 41.22 89.10 48.04 49.89 39.46 94.83 53.15 51.79 42.21 99.10 82.74 74.51 57.56
grid 92.13 45.75 48.84 27.66 89.18 34.89 41.21 19.21 85.19 24.32 34.76 18.22 98.78 62.24 58.44 41.69
hazelnut 59.97 28.87 37.82 30.38 99.36 95.16 89.80 76.43 98.54 77.39 70.42 45.97 98.97 88.00 81.77 73.39
leather 97.38 66.36 63.43 53.13 97.82 51.86 52.25 48.09 98.99 65.73 62.85 42.65 99.11 76.49 69.30 56.51
metal_nut 63.34 35.42 35.24 47.51 98.16 95.16 90.99 63.02 99.38 97.34 91.63 64.59 99.23 96.66 91.42 75.15
pill 36.18 8.93 12.79 13.62 95.80 75.90 70.31 31.73 98.96 92.51 85.35 50.04 98.11 79.63 72.54 56.73
screw 91.03 27.05 32.95 19.03 93.96 38.00 41.69 30.88 92.68 44.64 49.17 34.08 98.27 52.40 52.32 41.02
tile 91.77 80.02 77.19 56.27 97.37 88.79 82.05 66.30 92.98 79.59 73.52 55.08 99.38 96.24 89.90 75.50
toothbrush 55.94 24.87 35.54 12.75 95.17 55.21 53.95 28.83 98.31 68.60 66.14 29.67 96.97 54.84 53.19 27.91
transistor 55.81 19.36 24.38 32.58 97.68 89.68 84.18 46.98 98.20 83.97 75.84 44.22 98.80 84.32 77.02 55.57
wood 90.04 73.42 71.25 59.55 97.47 77.72 70.91 58.77 95.68 67.54 63.06 42.78 98.60 88.57 81.46 62.94
zipper 94.15 62.65 61.39 50.12 93.80 58.43 56.82 46.44 98.42 84.05 77.08 64.14 99.15 86.67 79.09 69.37
Average 76.41 40.99 43.91 35.61 94.91 67.06 65.09 45.49 96.30 69.77 66.99 45.77 98.01 77.77 72.81 57.62

### v2

@@ A.6 More qualitative and quantitative anomaly image generation results
Category DFMGAN [11] AnomalyDiffusion [17] Ours
@@
Figure 14: Qualitative results of our anomaly image generation results on MVTec AD. In the first row, from left to right, are the results for bottle, cable, and capsule categories. In the second row, from left to right, are the results for carpet, grid, and hazelnut categories.
@@
Figure 15: Qualitative results of our anomaly image generation results on MVTec AD. In the first row, from left to right, are the results for leather, metal_nut, and pill categories. In the second row, from left to right, are the results for screw, tile, and toothbrush categories.
Figure 16: Qualitative results of our anomaly image generation results on MVTec AD. In the first row, from left to right, are the results for transistor, wood, and zipper categories.
We provide further qualitative results of every category on the MVTec 3D AD dataset in Fig. 17. We report the anomaly image generation results of SeaS for varying types of anomalies. The first column represents the generated anomaly images, and the second column represents the corresponding generated masks.
Figure 17: Qualitative results of our anomaly image generation results on MVTec 3D AD. In the first row, from left to right, are the results for bagel, cable_gland, carrot, peach, and potato categories. In the second row, from left to right, are the results for cookie, dowel, foam, rope, and tire categories.
A.7 More qualitative and quantitative comparison results of supervised segmentation models trained on image-mask pairs generated by different anomaly generation methods
We provide further qualitative comparisons on downstream supervised segmentation trained by the generated images. The segmentation anomaly maps are shown in Fig. 22. There are fewer false positives (e.g., potato_combined) and fewer false negatives (e.g., bagel_contamination), when the BiSeNet V2 is trained on the image-mask pairs generated by our method.
Figure 18: Comparison results with the anomaly supervised segmentation model BiSeNet V2 on MVTec AD. In the figure, from top to bottom are the results for bottle, cable, capsule and carpet categories.
Figure 19: Comparison results with the anomaly supervised segmentation model BiSeNet V2 on MVTec AD. In the figure, from top to bottom are the results for grid, hazelnut, leather and metal_nut categories.
Figure 20: Comparison results with the anomaly supervised segmentation model BiSeNet V2 on MVTec AD. In the figure, from top to bottom are the results for pill, screw, tile and toothbrush categories.
Figure 21: Comparison results with the anomaly supervised segmentation model BiSeNet V2 on MVTec AD. In the figure, from top to bottom are the results for transistor, wood and zipper categories.
Figure 22: Qualitative supervised anomaly segmentation results with BiSeNet V2 on MVTec 3D AD.
@@
We report the detailed segmentation results of SeaS for each category on the MVTec AD datasets, compared with DFMGAN [11] and AnomalyDiffusion [17], which are presented from Tab. 22 to Tab. 27
Table 22: Comparison on supervised anomaly segmentation on BiSeNet V2.
Category DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU
bottle 89.34 64.67 62.78 44.71 99.00 88.02 80.53 68.25 99.46 93.43 85.59 75.86
cable 93.87 67.98 64.74 44.02 92.84 69.86 66.32 46.49 89.85 72.07 71.58 53.24
capsule 74.88 16.43 23.01 29.97 92.71 38.11 40.67 19.44 86.33 24.64 30.54 39.70
carpet 94.53 42.53 47.44 39.88 98.65 73.10 65.83 43.25 99.61 82.30 72.94 55.52
grid 96.86 24.40 37.40 29.93 80.59 8.08 16.79 14.26 99.36 37.91 42.50 39.80
hazelnut 99.87 96.75 90.07 71.68 97.71 63.34 59.87 43.12 97.82 78.55 73.09 68.47
leather 97.50 51.10 52.26 50.67 99.30 57.49 59.62 43.94 98.91 59.84 58.62 45.82
metal_nut 99.39 97.59 92.52 70.40 99.03 95.67 88.69 58.8 99.69 98.29 93.23 74.40
pill 97.09 83.98 79.26 36.39 99.44 93.16 86.62 41.18 98.31 76.97 68.00 55.43
screw 97.94 37.10 41.01 31.63 94.08 17.95 25.90 20.00 97.64 40.20 45.35 38.43
tile 99.65 97.08 91.16 75.94 97.79 85.58 78.28 60.46 99.67 97.29 91.48 75.75
toothbrush 97.70 51.32 54.05 23.38 98.43 49.64 54.08 26.53 97.15 46.09 49.02 28.56
transistor 84.31 45.34 46.07 30.00 98.85 85.27 77.95 49.83 96.75 69.52 66.11 57.24
wood 98.32 64.82 63.11 58.99 96.78 63.38 60.31 45.73 98.38 80.81 74.03 56.22
zipper 97.29 65.18 63.24 49.93 98.81 78.89 72.66 62.03 99.23 80.27 73.41 64.80
Average 94.57 60.42 60.54 45.83 96.27 64.5 62.27 42.89 97.21 69.21 66.37 55.28
@@
Category DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max
bottle 96.74 98.75 95.35 98.14 99.34 97.67 100.00 100.00 100.00
cable 79.47 85.00 74.13 95.37 96.71 92.91 94.61 96.39 89.83
capsule 85.51 95.16 89.82 84.06 95.01 89.74 88.81 96.92 89.21
carpet 91.42 96.29 88.89 90.55 96.41 90.32 98.16 99.31 97.56
grid 99.64 99.82 97.56 81.19 89.92 83.95 99.17 99.63 98.73
hazelnut 100.00 100.00 100.00 93.39 95.74 90.91 100.00 100.00 100.00
leather 98.31 99.23 95.24 100.00 100.00 100.00 95.83 98.38 95.93
metal_nut 97.37 99.16 94.66 99.01 99.66 97.71 100.00 100.00 100.00
pill 84.86 95.27 91.00 90.38 97.43 91.35 96.59 99.12 95.24
screw 74.95 85.50 80.72 58.18 75.32 81.25 77.24 89.55 80.60
tile 99.47 99.74 99.12 98.78 99.44 97.39 100.00 100.00 100.00
toothbrush 78.33 87.73 83.72 78.33 89.26 79.17 90.42 94.49 89.47
transistor 79.52 75.77 69.57 94.40 94.68 94.34 99.23 98.39 94.92
wood 98.87 99.46 97.67 90.48 94.12 93.33 100.00 100.00 100.00
zipper 98.97 99.64 97.56 98.89 99.62 97.56 100.00 100.00 100.00
Average 90.90 94.43 90.33 90.08 94.84 91.84 96.00 98.14 95.43
Table 24: Comparison on supervised anomaly segmentation on UPerNet.
Category DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU
bottle 87.94 56.89 56.56 45.41 99.54 93.01 85.94 75.31 99.28 91.73 84.53 78.73
cable 87.52 64.30 65.61 41.02 91.00 68.12 67.49 51.84 91.08 76.25 74.63 59.00
capsule 67.92 12.31 20.32 30.47 97.64 51.90 51.66 37.00 92.09 39.60 43.89 50.18
carpet 95.85 36.05 34.52 48.10 99.45 82.13 72.55 53.17 99.67 82.01 73.53 60.60
grid 97.49 29.67 36.15 31.37 94.22 28.97 38.50 32.93 99.18 44.94 48.28 44.21
hazelnut 99.36 79.76 71.10 72.90 97.77 70.48 67.93 54.47 99.54 81.84 75.48 73.30
leather 80.97 17.60 26.21 30.17 99.48 63.46 60.54 48.70 99.42 68.26 65.52 57.01
metal_nut 98.44 95.64 91.48 64.92 98.62 95.11 88.62 61.31 99.70 98.33 92.90 76.07
pill 97.58 83.74 80.02 42.33 99.33 95.04 88.77 49.18 98.59 81.16 74.26 62.62
screw 97.49 53.83 53.02 42.05 93.89 36.60 42.68 34.08 98.97 52.02 51.65 46.61
tile 99.79 97.29 91.11 77.46 94.70 73.34 67.79 58.54 99.67 95.89 90.71 77.89
toothbrush 97.42 51.09 59.23 28.33 97.52 60.67 59.46 33.98 98.50 63.62 63.07 42.09
transistor 82.07 36.31 39.48 27.44 94.26 73.68 69.50 53.64 93.88 70.37 68.12 56.98
wood 97.90 69.02 62.21 63.10 96.09 70.10 64.38 51.44 99.28 85.28 76.28 65.09
zipper 97.28 71.60 66.64 54.54 99.54 86.18 78.50 66.47 99.17 85.01 77.57 68.21
Average 92.33 57.01 56.91 46.64 96.87 69.92 66.95 50.80 97.87 74.42 70.70 61.24
@@
Category DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max
bottle 94.19 97.86 93.18 100.00 100.00 100.00 100.00 100.00 100.00
cable 85.64 90.03 80.33 95.58 97.06 92.56 94.40 96.38 92.44
capsule 81.04 94.26 87.01 96.00 98.77 95.48 94.43 98.44 92.21
carpet 96.72 98.58 93.75 98.68 99.53 98.36 99.94 99.97 99.20
grid 98.33 99.13 96.30 96.67 98.73 97.44 99.76 99.88 98.73
hazelnut 99.84 99.87 97.96 99.17 99.43 97.87 100.00 100.00 100.00
leather 79.91 90.70 81.75 100.00 100.00 100.00 100.00 100.00 100.00
metal_nut 98.30 99.38 97.71 98.65 99.62 98.41 99.72 99.91 99.21
pill 88.54 96.56 92.39 91.23 97.78 90.91 98.28 99.58 97.92
screw 89.01 94.54 88.24 85.06 93.87 85.33 93.47 97.07 90.45
tile 99.68 99.81 99.13 99.68 99.81 99.13 100.00 100.00 100.00
toothbrush 75.00 86.99 80.00 90.00 95.13 90.00 95.00 97.65 94.74
transistor 83.04 73.59 74.19 100.00 100.00 100.00 99.52 99.16 96.43
wood 93.36 95.60 95.45 98.62 99.49 97.62 99.87 99.94 98.82
zipper 98.48 99.51 98.14 100.00 100.00 100.00 100.00 100.00 100.00
Average 90.74 94.43 90.37 96.62 98.61 96.21 98.29 99.20 97.34
Table 26: Comparison on supervised anomaly segmentation on LFD.
Category DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU
bottle 90.41 61.51 58.49 40.19 98.71 89.64 81.55 67.10 99.28 92.65 84.86 73.82
cable 96.49 79.40 75.25 53.47 97.89 79.85 72.75 53.69 94.53 75.41 72.70 55.98
capsule 91.82 56.11 58.56 32.50 95.80 38.17 48.92 32.04 91.80 49.76 53.69 41.14
carpet 89.10 48.04 49.89 39.46 94.83 53.15 51.79 42.21 99.10 82.74 74.51 57.56
grid 89.18 34.89 41.21 19.21 85.19 24.32 34.76 18.22 98.78 62.24 58.44 41.69
hazelnut 99.36 95.16 89.80 76.43 98.54 77.39 70.42 45.97 98.97 88.00 81.77 73.39
leather 97.82 51.86 52.25 48.09 98.99 65.73 62.85 42.65 99.11 76.49 69.30 56.51
metal_nut 98.16 95.16 90.99 63.02 99.38 97.34 91.63 64.59 99.23 96.66 91.42 75.15
pill 95.80 75.90 70.31 31.73 98.96 92.51 85.35 50.04 98.11 79.63 72.54 56.73
screw 93.96 38.00 41.69 30.88 92.68 44.64 49.17 34.08 98.27 52.40 52.32 41.02
tile 97.37 88.79 82.05 66.30 92.98 79.59 73.52 55.08 99.38 96.24 89.90 75.50
toothbrush 95.17 55.21 53.95 28.83 98.31 68.60 66.14 29.67 96.97 54.84 53.19 27.91
transistor 97.68 89.68 84.18 46.98 98.20 83.97 75.84 44.22 98.80 84.32 77.02 55.57
wood 97.47 77.72 70.91 58.77 95.68 67.54 63.06 42.78 98.60 88.57 81.46 62.94
zipper 93.80 58.43 56.82 46.44 98.42 84.05 77.08 64.14 99.15 86.67 79.09 69.37
Average 94.91 67.06 65.09 45.49 96.30 69.77 66.99 45.77 98.01 77.77 72.81 57.62